OpenStudy (anonymous):

A cylinderical container is being filled with juice at a. Ra of pi cm^3/s. calculate the rate, after two seconds, at which the level of juice is rising, in cm/s. @satellite73 please help

OpenStudy (anonymous):

i guess it depends on the dimensions of the cylinder

OpenStudy (anonymous):

What o if i say the volume of cylinder is v=pi*r^2*h ?

OpenStudy (anonymous):

ok then \(V'=\pi r^2 h'\) since the radius if fixed you know \(V'=\pi\) so you can solve for \(h'\) although your answer will have an \(r\) in it

OpenStudy (anonymous):

Is the 'two seconds' matter?